### v1

# First-order CP phase transition in two-flavor QCD at $\theta=\pi$ under electromagnetic scale anomaly via a Nambu-Jona-Lasinio description

Abstract
We discuss the thermal CP phase transition in QCD at $\theta=\pi$ under a weak magnetic field background, where the electromagnetic scale anomaly gets significant. To explicitize, we work on a two-flavor Polyakov-loop Nambu-Jona-Lasinio model at $\theta=\pi$ in the mean field approximation, including the electromagnetic-scale anomaly term. We find that the thermal CP phase transition becomes first order and the strength of the first order gets more prominent as the magnetic field increases. The associated potential barrier is thermally created by the electromagnetic scale anomaly and gives rise to criticality due to the induced potential of a non-perturbative form $\sim\frac{|eB|^{3}}{f_{\pi}}\frac{|P|}{P^{2}+m_{0}^{2}}$, where $eB$ denotes the magnetic field strength; $P$ the CP order parameter, and $m_{0}$ the isospin-symmetric current-quark mass. The CP-broken deconfinement(-like) domain, $T^{\rm(CP)}_{c}>T^{\rm(dec)}_{\rm pc}$, gets wider as $eB$ increases.
I Introduction
Exploring CP violation is one of great importance to comprehend the cosmic history of the universe; e.g., pursuit of the origin of matter and anti-matter. This issue is essentially linked to a deep understanding of the topological feature of the non-Abelian gauge theories. In the Standard Model of particle physics, the so-called topological $\theta$ term in quantum chromodynamics (QCD) violates the CP symmetry unless $\theta=0$ or $\pi$. To the size of this $\theta$, the neutron electric dipole moment measurements have so far suggested $\theta$ to be $<10^{-11}$ Abel:2020pzs ; Liang:2023jfj . Thus the present-day strong CP violation has to be extremely small somehow with $\theta\sim 0$, not allowing for $\theta=\pi$.
The case might be quite different in the early hot universe: a local CP-odd domain may be created in hot QCD plasma due to the presence of the QCD sphaleron Manton:1983nd ; Klinkhamer:1984di , so that the QCD vacuum characterized by the strong CP phase $\theta$ and its fluctuation (in the spatial-homogeneous direction) gets significantly sizable Kharzeev:2007tn ; Kharzeev:2007jp ; Fukushima:2008xe within the QCD time scale McLerran:1990de ; Moore:1997im ; Moore:1999fs ; Bodeker:1999gx . Given this feasibility, QCD at $\theta=\pi$ is still well motivated to investigate in a viewpoint of not only the purely theoretical interest, but also even phenomenology or cosmology in the early universe.
The dynamics of the CP symmetry in Yang-Mills theories with or without quarks at $\theta=\pi$ has extensively been discussed so far. One of the key “benchmark”s is the so-called Dashen’s phenomenon Dashen:1970et , that suggested the spontaneous breaking of the CP symmetry at $\theta=\pi$ based on a chiral effective model approach. In earlier dates, instanton gas models and deconfinement models also developed this phenomenon to finite temperature, which involves the nature of the deconfinement phase transition Gross:1980br ; Weiss:1980rj : the CP symmetry and/or the center symmetry is spontaneously broken at lower temperatures, but is manifest at higher temperatures. This feature has been well captured in the large-$N_{c}$ limit of QCD with quarks Witten:1980sp ; Witten:1998uka , where the CP symmetry at $\theta=\pi$ is predicted to be spontaneously broken at low temperatures. Thus, CP symmetry $\theta=\pi$ is thought of also as a crucial ingredient in understanding the confinement.
Recently, the higher-form symmetry Gaiotto:2014kfa incorporated into the ’t Hooft anomaly matching tHooft:1979rat ; Frishman:1980dq has developed the study along this research line. It has been predicted that at $\theta=\pi$, either the CP or the center symmetry should be spontaneously broken unless the theory becomes gapless (i.e., deconfined): the theory cannot be in a gapped confined phase without breaking the CP symmetry Gaiotto:2017yup ; Cordova:2019bsd , namely, $T_{c}^{\rm(CP)}>T_{c}^{\rm(dec)}$.
For nonzero $\theta$, lattice simulations under the standard base of the MonteCarlo method unfortunately suffers from infamous sign problem. To avoid the sign problem, interesting methods for nonzero $\theta$ have so far been proposed. In particular, the so-called subvolume method has been applied to four-dimensional $SU(2)$ Yang-Mills theory, which suggests that the CP symmetry at $\theta=\pi$ is broken at zero temperature and is restored at higher temperatures Kitano:2020mfk ; Kitano:2021jho ; Yamada:2024pjy ; Yamada:2024vsk . Furthermore, very recently, another technique based on the analytic continuation from the imaginary $\theta$ to real $\theta$ has been applied to the lattice simulation and also observed the same phenomenon Hirasawa:2024fjt .
In a realistic setup including quarks, where $\theta$ is transformed by the axial rotation to be able to have dependence on the number of quark flavors $N_{f}$ as $\theta\to\theta_{N_{f}}=\theta/N_{f}$ with identical masses, the analytic continuation method can still be utilized to get expectation values for physical observables DelDebbio:2002xa ; DElia:2003zne ; DelDebbio:2006sbu ; Giusti:2007tu ; Izubuchi:2007rmy ; Vicari:2008jw ; DElia:2013uaf ; Hirasawa:2024vns . However, those calculations have been done away from the physical point. The subvolume method proposed in Kitano:2021jho is also applicable even in the case with quarks, but it is still hard to reach physical results. Another idea based on digital and analog quantum simulations and calculations has been adopted to the case with nonzero $\theta$ using tensor networks, which have been applied only for toy models, but not for the realistic setup like four dimensional QCD at the physical point Byrnes:2002gj ; Funcke:2019zna ; Kuramashi:2019cgs ; Chakraborty:2020uhf ; Honda:2021aum ; Honda:2022hyu . Thus, reaching realistic QCD with quarks is still challenging for lattice simulations, so the spontaneous CP symmetry breaking and its restoration still involves lots of unknowns.
On the other hand, QCD at $\theta=\pi$ has so far been explored also by various chiral effective models Dashen:1970et ; Witten:1980sp ; Pisarski:1996ne ; Creutz:2003xu ; Mizher:2008hf ; Boer:2008ct ; Creutz:2009kx ; Boomsma:2009eh ; Sakai:2011gs ; Chatterjee:2011yz ; Sasaki:2011cj ; Sasaki:2013ewa ; Aoki:2014moa ; Mameda:2014cxa ; Verbaarschot:2014upa ; Bai:2023cqj ; Huang:2024nbd . The field of target physics is nowadays getting broader, which has been originated from the Dashen phenomenon up until the recent interest in gravitational wave productions based on the expected CP phase transition of the first-order or the domain wall collapse of the axionlike particle at $\theta=\pi$. Those studies have provided benchmarks toward understanding of QCD at $\theta=\pi$, prior to the future development in lattice simulations.
In light of the early universe (and/or heavy ion collision experiments), however, the CP phase transition might be more involved: a strong electromagnetic field could be generated via cosmological first-order phase transitions Vachaspati:1991nm ; Enqvist:1993np ; Grasso:1997nx ; Grasso:2000wj ; Ellis:2019tjf ; Zhang:2019vsb ; Di:2020kbw ; Yang:2021uid ; Saga:2023afb ; Liu:2024mdo ; Di:2024gsl , or inflationary scenarios Turner:1987vd ; Garretson:1992vt ; Anber:2006xt ; Domcke:2019mnd ; Domcke:2019qmm ; Patel:2019isj ; Domcke:2020zez ; Shtanov:2020gjp ; Okano:2020uyr ; Cado:2021bia ; Kushwaha:2021csq ; Gorbar:2021rlt ; Gorbar:2021zlr ; Gorbar:2021ajq ; Fujita:2022fwc ; Cado:2022pxk ; Domcke:2022kfs ; Cado:2023gan ; Murai:2023gkv ; Bastero-Gil:2023htv ; Bastero-Gil:2023mxm ; Otsuka:2024mdg ; vonEckardstein:2024tix ; Sharma:2024nfu . The strength of the produced electromagnetic field would be redshifted to be smaller and smaller toward a later time epoch including the timing when the universe undergoes the QCD phase transition. With the active QCD sphaleron taken into account McLerran:1990de ; Moore:1997im ; Moore:1999fs ; Bodeker:1999gx , thereby, the CP phase transition at $\theta=\pi$ can be under such a redshifed background electromagnetic field.
Although the electromagnetic field does not break the CP symmetry additionally, it does explicitly break the chiral and $U(1)$ axial symmetries because of the different charges for quarks. This breaking can cause the so-called magnetic catalysis or inverse magnetic catalysis Bali:2011qj , where the constituent dynamical quark masses and the quark condensates are amplified or suppressed by the applied magnetic field. At $\theta=\pi$ (i.e., $\theta_{N_{f}}=\pi/N_{f}$), the scalar quark condensate will be interchanged with the pseudoscalar quark condensate via the axial rotation, where the latter plays the role of the order parameter for the CP symmetry breaking as well as the chiral order parameter. Therefore, the CP symmetry breaking in the same way as in the chiral symmetry breaking is expected to be enhanced or suppressed in the presence of a magnetic field. In fact, this kind of phenomenon has been observed in the chiral effective model approaches Chatterjee:2014csa ; Bandyopadhyay:2019pml ; Carlomagno:2025ayh .
In a view of a cosmic magnetic field background at the QCD epoch as aforementioned, the magnetic field strength might be redshifted to be as small as or smaller than the dynamical quark mass scale $m_{\rm dyn}\sim 300-400$ MeV: $eB\lesssim m_{\rm dyn}^{2}$. In that case, it has recently been clarified that the electromagnetic-scale anomaly effect is not negligible in a weak magnetic field and is shown to have important implications for the chiral phase transition properties regarding the quark mass dependence in QCD with $\theta=0$ Kawaguchi:2021nsa ; YuanyuanWang:2022nds ; YuanyuanWang:2024zwf . This scale anomaly effect has never been examined along with nonzero $\theta$.
In this paper, we discuss the effect of the electromagnetic scale anomaly by explicitly working on a two-flavor NJL model at $\theta=\pi$ with the Polyakov loop extension (PNJL model) in the mean field approximation (MFA). We observe that the thermomagnetic correction part in the electromagnetic scale anomaly term develops a potential barrier around the origin of the direction of the CP order parameter. It turns out that this induced potential term is of nonperturbative form, which cannot be described by the Ginzburg-Landau description.
The phase transitions along the chiral and Polyakov loop directions are smooth crossover. The deconfinement crossover goes not much different from the case without the electromagnetic scale anomaly. The CP-broken deconfinement(-like) domain is thus emergent: $T^{\rm(CP)}_{c}>T^{\rm(dec)}_{\rm pc}$, and it gets enlarged with a larger size of $eB$.
II PNJL model with electromagnetic scale anomaly
We begin by introducing the two-flavor PNJL model with nonzero $\theta$ under a constant external magnetic field. The Lagrangian of this model is as follows:
$\mathcal{L}=$ $\bar{\psi}\left(i\gamma^{\mu}D_{\mu}-m_{0}\cdot{\bf 1}_{2\times 2}\right)\psi+\frac{g_{s}}{2}\sum_{a=0}^{3}\left[\left(\bar{\psi}\tau_{a}\psi\right)^{2}+\left(\bar{\psi}i\gamma_{5}\tau_{a}\psi\right)^{2}\right]$ (1)
$+g_{d}\left\{e^{i\theta}\operatorname{det}\left[\bar{\psi}\left(1+\gamma_{5}\right)\psi\right]+e^{-i\theta}\operatorname{det}\left[\bar{\psi}\left(1-\gamma_{5}\right)\psi\right]\right\}-\mathcal{U}(\Phi,\bar{\Phi};T),$
where $\psi=(u,d)^{T}$ denotes the two-flavor quark field and the isospin-symmetric current-quark mass $m_{0}$. The covariant derivative $D_{\mu}=\partial_{\mu}-iqA^{EM}_{\mu}-iA_{\mu}$ contains the couplings of quarks to the gluon field $A_{\mu}$, and the external electromagnetic field $A^{EM}_{\mu}$, with its electromagnetic charge matrix given by $q=e\cdot\operatorname{diag}\{2/3,-1/3\}$. For simplicity, the magnetic field $B$ is set along the $z$-axis and is embedded in the electromagnetic gauge field of the form $A^{EM}_{\mu}=(0,0,Bx,0)$. The gluon field $A_{\mu}$ includes only the temporal component $A_{\mu}=\delta_{\mu}^{0}A_{0}=-iA_{4}$, where $A_{4}$ denotes the field in Euclidean space. The four-fermion interaction with the coupling $g_{s}$ has the full chiral $U(2)_{L}\times U(2)_{R}$ invariance, and $\tau_{a}(a=0,\ldots,3)$ denotes the $U(2)$ generators including $\tau_{0}={\bf 1}_{2\times 2}$. The term with the coupling $g_{d}$ is called the Kobayashi-Maskawa-’t Hooft (KMT) determinant term Kobayashi:1970ji ; Kobayashi:1971qz ; tHooft:1976rip ; tHooft:1976snw and preserves the $SU(2)_{L}\times SU(2)_{R}\times U(1)_{V}$ invariance, but breaks the $U(1)$ axial symmetry.
[…]
$\mathcal{U}(\Phi,\bar{\Phi};T)=T^{4}\left[-\frac{b_{2}(T)}{2}\bar{\Phi}\Phi-\frac{b_{3}}{6}\left(\bar{\Phi}^{3}+\Phi^{3}\right)+\frac{b_{4}}{4}(\bar{\Phi}\Phi)^{2}\right],$ (2)
where
$b_{2}(T)=a_{0}+a_{1}\left(\frac{T_{0}}{T}\right)+a_{2}\left(\frac{T_{0}}{T}\right)^{2}+a_{3}\left(\frac{T_{0}}{T}\right)^{3}.$ (3)
The parameter settings in the Polyakov potential are fixed by the pure Yang-Mills lattice simulation with the following values Meisinger:2001cq : $a_{0}=6.75$, $a_{1}=-1.95$, $a_{2}=2.625$, $a_{3}=-7.44$, $b_{3}=0.75$ and $b_{4}=7.5$. $T_{0}=0.27\,\mathrm{GeV}$ denotes the critical temperature of the deconfinement phase transition in pure Yang-Mills theory.
We introduce the three-dimensional momentum cutoff $\Lambda$ to regularize the PNJL model in the MFA. We refer to empirical hadron observables in the isospin symmetric limit at $eB=T=0$ Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs : the pion mass $m_{\pi}=0.1402\,\mathrm{GeV}$, the pion decay constant $f_{\pi}=0.0926\,\mathrm{GeV}$, and quark condensate $\langle\bar{u}u\rangle=\langle\bar{d}d\rangle=(-0.2415\,\mathrm{GeV})^{3}$. Then the model parameters can be determined as $m=0.006\,\mathrm{GeV}$, $\Lambda=0.59\,\mathrm{GeV}$, and for $g_{s}=2(1-c)G_{0}$ with $g_{d}=2cG_{0}$ and $c=0.2$ Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs , we have $G_{0}\Lambda^{2}=2.435$.
We utilize the imaginary time formalism and the Landau level decomposition to evaluate the thermomagnetic contributions to the quark condensate taking the following replacements:
$p_{0}$ $\leftrightarrow i\omega_{\mathbf{k}}=i(2\mathbf{k}+1)\pi T$ (4)
$\int\frac{d^{4}p}{(2\pi)^{4}}$ $\leftrightarrow iT\sum_{\mathbf{k}=-\infty}^{\infty}\int\frac{d^{3}p}{(2\pi)^{3}}$
$\leftrightarrow iT\sum_{\mathbf{k}=-\infty}^{\infty}\sum_{n=0}^{\infty}\alpha_{n}\frac{\left|q_{f}B\right|}{4\pi}\int_{-\infty}^{\infty}\frac{dp_{z}}{2\pi}f_{\Lambda}\left(p_{z},n\right).$
$\omega_{k}$ represents the Matsubara frequency with $k$ as an integer, while $n$ denotes the Landau levels. The factor $\alpha_{n}=2-\delta_{n,0}$ refers to the spin degeneracy for the Landau levels. By applying these replacements, the chiral phase transition can be directly extended from the vacuum to the case of $T\neq 0$ and $eB\neq 0$. In addition we introduce a soft-cutoff regulator function Frasca:2011zn ,
$f_{\Lambda}\left(p_{z},n\right)=\frac{\Lambda^{10}}{\Lambda^{10}+\left(p_{z}^{2}+2n\left|q_{f}B\right|\right)^{5}}\,.$ (5)
It is convenient to transfer the $\theta$ dependence attached on the $g_{d}$ term in Eq.(1) from the $g_{d}$ term to the mass term by a $U(1)$ axial rotation of $\psi$ as $\psi\rightarrow e^{-i\gamma_{5}\frac{\theta}{4}}\psi\equiv\psi^{\prime}$. Then the Lagrangian in Eq.(1) can be rewritten as
$\mathcal{L}\rightarrow\mathcal{L}^{\prime}=$ $\bar{\psi^{\prime}}\left(i\gamma^{\mu}D_{\mu}-\mathbf{m}(\theta)\right)\psi^{\prime}+\frac{g_{s}}{2}\sum_{a=0}^{3}\left[\left(\bar{\psi^{\prime}}\tau_{a}\psi^{\prime}\right)^{2}+\left(\bar{\psi^{\prime}}i\gamma_{5}\tau_{a}\psi^{\prime}\right)^{2}\right]$ (6)
$+g_{d}\left\{\operatorname{det}\left[\bar{\psi^{\prime}}\left(1+\gamma_{5}\right)\psi^{\prime}\right]+\operatorname{det}\left[\bar{\psi^{\prime}}\left(1-\gamma_{5}\right)\psi^{\prime}\right]\right\}-\mathcal{U}(\Phi,\bar{\Phi};T),$
where
$\mathbf{m}(\theta)$ $=m_{0}\left[\cos\frac{\theta}{2}+i\gamma_{5}\sin\frac{\theta}{2}\right]\,,$
$=m_{0}\left[\cos\theta_{N_{f}=2}+i\gamma_{5}\sin\theta_{N_{f}=2}\right]\,,$ (7)
and all the $\theta$ dependence has been now transformed into the complex mass $\mathbf{m}(\theta)$. The scalar and pseudoscalar bilinears in the original base are related to those in the prime base:
$\left(\bar{\psi}\psi\right)$ $=\left(\bar{\psi}^{\prime}\psi^{\prime}\right)\cos\frac{\theta}{2}+\left(\bar{\psi}^{\prime}i\gamma_{5}\psi_{i}^{\prime}\right)\sin\frac{\theta}{2}\,,$ (8)
$\left(\bar{\psi}i\gamma_{5}\psi\right)$ $=-\left(\bar{\psi}^{\prime}\psi^{\prime}\right)\sin\frac{\theta}{2}+\left(\bar{\psi}^{\prime}i\gamma_{5}\psi^{\prime}\right)\cos\frac{\theta}{2}.$
We work in the MFA so that the isospin-singlet scalar-quark bilinear $\bar{\psi}\psi=\bar{u}u+\bar{d}d$ and its CP partner $\bar{\psi}i\gamma_{5}\psi=\bar{u}i\gamma_{5}u+\bar{d}i\gamma_{5}d$ can be expanded around the mean fields $S_{q}^{\prime}=\left\langle\bar{q}^{\prime}q^{\prime}\right\rangle(q=u,d)$, and $P_{q}^{\prime}=\left\langle\bar{q}^{\prime}i\gamma_{5}q^{\prime}\right\rangle(q=u,d)$. The MF values for other isotriplet quark-bilinears are set to zero. Thus, in the MFA, the Lagrangian in Eq.(6) takes the form
$\mathcal{L}_{\mathrm{MFA}}=$ $\bar{\psi}^{\prime}\left(i\gamma^{\mu}D_{\mu}-\mathbf{M}(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime})\right)\psi^{\prime}-\frac{g_{s}}{2}\left[\left(S_{u}^{\prime}+S_{d}^{\prime}\right)^{2}+\left(P_{u}^{\prime}+P_{d}^{\prime}\right)^{2}\right]$ (9)
$-2g_{d}\left(S_{u}^{\prime}S_{d}^{\prime}-P_{u}^{\prime}P_{d}^{\prime}\right)-\mathcal{U}(\Phi,\bar{\Phi};T),$
where
$\mathbf{M}(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime})=\operatorname{diag}\{M_{u}(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime}),M_{d}(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime})\}\,,$ (10)
and
$M_{u}(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime})$ $=\alpha_{u}\left(\theta,S_{u}^{\prime},S_{d}^{\prime}\right)+i\gamma_{5}\beta_{u}\left(\theta,P_{u}^{\prime},P_{d}^{\prime}\right)\,,$ (11)
$\alpha_{u}\left(\theta,S_{u}^{\prime},S_{d}^{\prime}\right)$ $=m_{0}\cos\frac{\theta}{2}-g_{s}\left(S_{u}^{\prime}+S_{d}^{\prime}\right)-2g_{d}S_{d}^{\prime},$
$\equiv m_{0}\cos\frac{\theta}{2}+\sigma_{u}^{\prime}\,,$
$\beta_{u}\left(\theta,P_{u}^{\prime},P_{d}^{\prime}\right)$ $=m_{0}\sin\frac{\theta}{2}-g_{s}\left(P_{u}^{\prime}+P_{d}^{\prime}\right)+2g_{d}P_{d}^{\prime},\,,$
$\equiv m_{0}\sin\frac{\theta}{2}+\eta_{u}^{\prime}\,,$
$M_{d}(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime})$ $=\alpha_{d}\left(\theta,S_{u}^{\prime},S_{d}^{\prime}\right)+i\gamma_{5}\beta_{d}\left(\theta,P_{u}^{\prime},P_{d}^{\prime}\right),$
$\alpha_{d}\left(\theta,S_{u}^{\prime},S_{d}^{\prime}\right)$ $=m_{0}\cos\frac{\theta}{2}-g_{s}\left(S_{u}^{\prime}+S_{d}^{\prime}\right)-2g_{d}S_{u}^{\prime},$
$=m_{0}\cos\frac{\theta}{2}+\sigma_{d}^{\prime}\,,$
$\beta_{d}\left(\theta,P_{u}^{\prime},P_{d}^{\prime}\right)$ $=m_{0}\sin\frac{\theta}{2}-g_{s}\left(P_{u}^{\prime}+P_{d}^{\prime}\right)+2g_{d}P_{u}^{\prime},,$
$\equiv m_{0}\sin\frac{\theta}{2}+\eta_{d}^{\prime}\,.$
The thermomagnetic potential in the MFA is obtained by integrating out the quarks as follows:
$\Omega\left(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},\Phi,\bar{\Phi},T,eB\right)$ $=\frac{g_{s}}{2}\left[\left(S_{u}^{\prime}+S_{d}^{\prime}\right)^{2}+\left(P_{u}^{\prime}+P_{d}^{\prime}\right)^{2}\right]+2g_{d}\left(S_{u}^{\prime}S_{d}^{\prime}-P_{u}^{\prime}P_{d}^{\prime}\right)$ (12)
$-\sum_{f=u,d}\sum_{n=0}^{\infty}\alpha_{n}\int\frac{dp_{z}}{2\pi}\frac{\left|q_{f}B\right|}{2\pi}f_{\Lambda}\left(p_{z},n\right)\left[3E_{f,n}\right.$
$+T\ln\left(1+3\Phi e^{-\frac{E_{f,n}}{T}}+3\bar{\Phi}e^{-2\frac{E_{f,n}}{T}}+e^{-3\frac{E_{f,n}}{T}}\right)$
$\left.+T\ln\left(1+3\bar{\Phi}e^{-\frac{E_{f,n}}{T}}+3\Phi e^{-2\frac{E_{f,n}}{T}}+e^{-3\frac{E_{f,n}}{T}}\right)\right]+\mathcal{U}(\Phi,\bar{\Phi};T),$
where
$M^{2}_{f}$ $=\alpha^{2}_{f}+\beta^{2}_{f}\,,$
$E_{f,n}$ $=\sqrt{M^{2}_{f}+2n\left|q_{f}B\right|+p^{2}_{z}}\,,$ (13)
for $f=u$ and $d$.
Besides, we incorporate the electromagnetic scale anomaly into the PNJL model Kawaguchi:2021nsa ; YuanyuanWang:2022nds ; YuanyuanWang:2024zwf :
$V_{\mathrm{eff}}^{(\mathrm{Tad})}=-\frac{\varphi}{f_{\varphi}}T_{\mu}^{\mu}\,,$ (14)
where
$\varphi=\sqrt{\sum_{f=u,d}\left[(\sigma_{f}^{\prime})^{2}+(\eta_{f}^{\prime})^{2}\right]}\,,$ (15)
and $f_{\varphi}$ stands for the vacuum expectation value of this $\varphi$ at $T=eB=0$. What we are concerned about is in a weak field limit, $eB\ll m_{\rm dyn}^{2}$, as noted in the Introduction #1#1#1 When $eB$ exceeds the threshold strength $\gtrsim m_{\rm dyn}^{2}$, the electromagnetic scale anomaly gets started to suppress. This is due to the fact that above the threshold, the dynamics of quarks will be dominated by the lowest Landau level states, which are longitudinally polarized along the magnetic field direction. Since the electromagnetic scale anomaly develops along the transverse direction, its effect becomes mostly remarkable in the weak magnetic regime Kawaguchi:2021nsa .. Then the trace anomaly $T_{\mu}^{\mu}$ in Eq.(14) takes the form YuanyuanWang:2024zwf
$T_{\mu}^{\mu}=\frac{\beta(e)}{e^{3}}|eB|^{2}+\frac{1}{2}N_{c}\sum_{f=u,d,s}\sum_{n=0}^{\infty}\alpha_{n}\frac{q_{f}^{2}}{e^{2}}M^{2}_{f}\frac{\left|q_{f}B\right|}{4\pi}\int_{-\infty}^{\infty}\frac{dp_{z}}{2\pi}F\left(T,eB,M_{f}\right)|eB|^{2}.$ (16)
The first term $\beta(e)$ in Eq.(16) denotes the beta-function coefficient of the electromagnetic coupling $e$ at the one-loop level, which is evaluated as
$\beta(e)=\frac{1}{(4\pi)^{2}}\frac{4N_{c}}{3}\sum_{f}e\,q_{f}^{2}\,,$ (17)
and represents the electromagnetic scale anomaly contribution in the magnetized vacuum without $T$. The second term in Eq.(16) is the thermomagnetic part including the Fermi-Dirac thermal distribution function as #2#2#2Higher-order contributions in powers of $eB$ are potentially included in the full quark mass $M_{f}$ in the second term of Eq.(16) , while other possible higher order terms in $eB$ have been disregarded in Eq.(16). This may be a crude truncation. In Ghosh:2019kmf a nonperturbative evaluation of the photon polarization function coupled to the chiral singlet meson field has been made attempted. This could it possible to make a straightforward computation of the photon polarization in the weak magnetic field. The reliability of the present truncation prescription could be evaluated in such a way, which is to be left in the future work.
$F\left(T,eB,M_{f}\right)=\frac{1}{E^{5}_{f,n}}\cdot\frac{-2}{\exp\left(\frac{E_{f,n}}{T}\right)+1},\qquad{\rm with}\qquad$ $E^{2}_{f,n}=p_{z}^{2}+2n\left|q_{f}B\right|+M_{f}^{2},$ (18)
The total thermomagnetic potential in the MFA at nonzero $\theta$ containing the Polyakov loop potential and the tadpole potential is thus given as the sum of two contributions in Eqs.(12) and (14) with Eq.(16),
$\Omega_{tot}\left(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},\Phi,\bar{\Phi},T,eB\right)=\Omega\left(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},\Phi,\bar{\Phi},T,eB\right)+V_{\text{eff}}^{(\mathrm{Tad})}\left(\theta,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},T,eB\right).$ (19)
When $T$ and $eB$ are given with $\theta=\pi$, the order parameters of the chiral, CP, and deconfinement(-like) phase transitions can be determined by the following stationary conditions:
$\frac{\partial\Omega_{tot}\left(\theta=\pi,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},\Phi,\bar{\Phi},T,eB\right)}{\partial S_{f}^{\prime}}=0,\quad\frac{\partial\Omega_{tot}\left(\theta=\pi,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},\Phi,\bar{\Phi},T,eB\right)}{\partial P_{f}^{\prime}}=0,$ (20)
$\frac{\partial\Omega_{tot}\left(\theta=\pi,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},\Phi,\bar{\Phi},T,eB\right)}{\partial\Phi}=0,\quad\frac{\partial\Omega_{tot}\left(\theta=\pi,S_{u}^{\prime},S_{d}^{\prime},P_{u}^{\prime},P_{d}^{\prime},\Phi,\bar{\Phi},T,eB\right)}{\partial\bar{\Phi}}=0\,.$
III CP phase transition at $\theta=\pi$
Figure 1: Plots of $\frac{P_{u}}{P_{u}(T=eB=0)}=\frac{\left\langle\bar{u}i\gamma_{5}u\right\rangle}{\left\langle\bar{u}i\gamma_{5}u\right\rangle(T=eB=0)}$ versus temperature for varying magnetic fields, at $\theta=\pi$. Left panel shows the case without the tadpole potential term in Eq.(14) with Eq.(16), while the right panel with the tadpole contribution.
In Fig. 1, we plot the $T$ dependence of $P_{u}=\left\langle\bar{u}i\gamma_{5}u\right\rangle$ normalized to its vacuum value at $T=eB=0$ for various magnetic field strengths, with (right panel) and without (left panel) the tadpole potential contribution. Since the down-sector part $P_{d}$ shows qualitatively the same behavior against $T$ as what $P_{u}$ does, we will not explicitly display it.
In the case with $eB\neq 0$, but without the tadpole contributions (see the left panel in Fig. 1), the order parameter for both the chiral and CP symmetries, $P$, undergoes the normal second-order phase transition at the criticality (around $\sim 300$ MeV). This phase transition is fairly insensitive to the presence of the weak $eB$ (See comparison with the curve for $eB=0$#3#3#3 The critical point of the CP phase transition at $eB=0$ observed in the current analysis is slightly different from what has been seen in the literature Sakai:2011gs ; Huang:2024nbd based on the same PNJL model with the same set of the input parameters. This is because of the different regularization scheme: the smooth cutoff regulator is used as done in Eq.(5) or not., although the overall size of $P$ gets slightly amplified by $eB\neq 0$ due to the magnetic catalysis for the chiral symmetry breaking, as has been observed in the literature Chatterjee:2014csa ; Bandyopadhyay:2019pml .
Interestingly enough, with the tadpole term contribution (the right panel of Fig. 1), the CP phase transition becomes first order and the criticality is shifted towards higher temperatures as $eB$ increases. This first order nature emerges due to the presence of the potential barrier generated by the thermomagnetic correction part in the tadpole term, Eq.(14) with Eq.(16). Figure 2 visualizes the tadpole term contribution to the total potential projected along the $P$-direction, where we have chosen $eB=0.02\,{\rm GeV}^{2}$, various temperatures around the criticality. For simplicity, we have taken the isospin symmetric limit: $P^{\prime}_{u}=P^{\prime}_{d}=0$, i.e. $\beta_{u}=\beta_{d}=\beta=m_{0}$, $S^{\prime}_{u}=S^{\prime}_{d}=S^{\prime}=-\alpha/(2(g_{s}+g_{d}))$, $q_{u}=q_{d}=(2/3)e$. One can see from the figure that the tadpole contribution generates the barrier around the origin of the $P$-direction.
Figure 2: The deformation of the thermomagnetic potential projected onto the $\alpha\propto P$ direction, for the case with the tadpole potential term, in comparison with the case without the tadpole term, at $\theta=\pi$, $eB=0.02\,{\rm GeV}^{2}$, for various values of $T$ around the criticality ($T=T_{c}\simeq 305.7$ MeV). For simplicity, we have taken the isospin symmetric limit, where $P^{\prime}_{u}=P^{\prime}_{d}=0$ (i.e. $\beta_{u}=\beta_{d}=\beta=m_{0}$), $S^{\prime}_{u}=S^{\prime}_{d}=S^{\prime}=-\alpha/(2(g_{s}+g_{d}))$, $q_{u}=q_{d}=(2/3)e$, and $\Phi=\bar{\Phi}\simeq 0.7$ around $T=300$ MeV. Potential values have been adjusted by dividing by a factor of $(0.1\,{\rm GeV})^{4}$, just for convenience.
To clarify the barrier generation mechanism, in the left panel of Fig. 3 we plot the tadpole term contributions at $\theta=\pi$ as a function of $\alpha_{u}$ in Eq.(11), for $T=200$ MeV and 300 MeV with $eB=0.02\,{\rm GeV}^{2}$ fixed. The potential has been normalized as $V_{\rm eff}^{\rm Tad}(0)=0$. Again, for simplicity, we have taken the isospin symmetric limit as above. A spike structure is observed indeed in the thermomagnetic part (blue curves), which is not substantially affected even in the total tadpole term (black curves) including the vacuum term at $T=0$ (red curve). This sort of a peak structure is not seen when $\theta=0$ along the $\alpha$ direction (right panel of Fig. 3), where the isospin symmetric limit reads $P_{u}=P_{d}=0$, i.e., $\beta_{u}=\beta_{d}=\beta=0$, and $S_{u}=S_{d}=S=-(\alpha-m_{0})/(2(g_{s}+g_{d}))$.
Figure 3: (Left panel): the breakdown of the tadpole-potential term-contributions in Eq.(14), as a function of the CP order parameter $\alpha(\theta=\pi)=-2P(g_{s}+g_{d})$ normalized by the vacuum value at $T=eB=0$. The plots are split into the full contribution (black curve), the vacuum part at $T=0$ (red curve), and the thermomagnetic part (blue curve), for $T=200$ MeV and 300 MeV, with $eB=0.02\,{\rm GeV}^{2}$ fixed. The isospin symmetric limit has been applied as described in the main text. (Right panel): the same for $\alpha(\theta=0)=-2S(g_{s}+g_{d})$. The potential values in both panels have been adjusted by multiplying by some factor as in Fig. 2.
This discrepancy can be understood in an analytic way. First of all, consider only the thermomagnetic term in Eq.(14) (i.e., the second term in Eq.(16)) at $\theta=\pi$, and take the isospin symmetric limit as above. Then the corresponding tadpole potential term goes like
$V_{\rm eff}^{\rm(Tad)}(\alpha)\Bigg{|}_{\theta=\pi}\approx\frac{|eB|^{3}}{f_{0}}\frac{(2/3)^{2}}{8\pi^{2}}\sum_{n=0}^{\infty}\alpha_{n}\frac{|\alpha|}{{\bf M}_{n}^{2}(\alpha)}{\cal I}(T/{\bf M}_{n}(\alpha))\,,$ (21)
where $\alpha$ at $\theta=\pi$ is related to $P$ in the original base as $\alpha=-2(g_{s}+g_{d})P$, and
${\bf M}^{2}_{n}(\alpha)$ $=\alpha^{2}+\frac{4n}{3}|eB|+m_{0}^{2}\,,$
${\cal I}(T/{\bf M}_{n}(\alpha))$ $=\int_{0}^{\infty}\frac{dz}{(z^{2}+1)^{5/2}}\frac{1}{1+e^{\frac{\sqrt{z^{2}+1}}{T/{\bf M}_{n}(\alpha)}}}\,.$ (22)
It is manifest from Eq.(21) that this thermomagnetic contribution becomes significant at higher $T$ ($T\gtrsim{\bf M}_{n}(\alpha)$): otherwise Boltzmann suppressed. At around the criticality $T\sim 300$ MeV, the dynamical quark mass $(\alpha)$ has already dropped down to be smaller than the current quark mass $m_{0}$ (see Fig. 1), at which moment with $eB={\cal O}(f_{\pi}^{2})$, only the lowest Landau level essentially contributes. Thereby, Eq.(21) can further be approximated to be
$V_{\rm eff}^{\rm(Tad)}(\alpha)\Bigg{|}_{\theta=\pi}^{T\sim 300\,{\rm MeV};eB={\cal O}(f_{\pi}^{2})}\approx\frac{|eB|^{3}}{f_{0}}\frac{(2/3)^{2}}{8\pi^{2}}\frac{|\alpha|}{\alpha^{2}+m_{0}^{2}}{\cal I}(T/{\bf M}_{0}(\alpha))\,.$ (23)
Now the integral ${\cal I}_{\pi}$ is almost constant in $\alpha$, so that this tadpole term scales with $\alpha$ and $eB$ as
$V_{\rm eff}^{\rm(Tad)}(\alpha)\Bigg{|}_{\theta=\pi}^{T\sim 300\,{\rm MeV};eB={\cal O}(f_{\pi}^{2})}\propto\frac{|eB|^{3}}{f_{0}}\frac{|\alpha|}{\alpha^{2}+m_{0}^{2}}\,.$ (24)
This has a peak at $\alpha=m_{0}$, hence a potential barrier is built there, and its magnitude gets greater as $eB$ increases. The peak structure appears because the method of reducing the quark-thermal loop integral from three dimensions to just one dimension (along the $p_{z}$ direction) works well. In this process, the spatial momentum integral is replaced by the area related to the magnetic field ($eB$) and a sum over Landau levels. Near the critical point, the lowest Landau level dominates the contribution to the tadpole term. This allows the dependence on $\alpha$ to be determined mainly by the scaling of the one-dimensional loop momentum $p_{z}$, as shown in Eq.(21). If the scaling of the other spatial momenta ($p_{x},p_{y}$) or alternatively higher Landau levels were considered, it would introduce additional $\alpha$-dependence, potentially disrupting the barrier generation structure.
When this tadpole term is combined with the non-anomalous potential part ($\Omega$ in Eq.(19)), higher barrier tends to shift the criticality to be higher $T$. This trend is induced due to the cubic power scaling $\sim|eB|^{3}$ in Eq.(24). This magnetic catalysis has indeed been observed in the right panel of Fig. 1. Of particular interest is also to note that the $\alpha$ dependence in Eq.(24) is totally nonperturbative: beyond a simple-minded Ginzburg-Landau description.
On the other hand, at $\theta=0$, the thermomagnetic part in Eq.(14) is evaluated as
$V_{\rm eff}^{\rm(Tad)}(\alpha)\Bigg{|}_{\theta=0}\approx\frac{|eB|^{3}}{f_{0}}\frac{(2/3)^{2}}{8\pi^{2}}\sum_{n=0}^{\infty}\alpha_{n}\frac{|\alpha-m_{0}|}{{\bf M^{\prime}}_{n}^{2}(\alpha)}{\cal I}(T/{\bf M^{\prime}}_{n}(\alpha))\,.$ (25)
In this case $\alpha$ at $\theta=0$ is related to $S$ in the original base as $\alpha=-2(g_{s}+g_{d})S$, and
${\bf M^{\prime}}^{2}_{n}(\alpha)$ $=\alpha^{2}+\frac{4n}{3}|eB|\,.$ (26)
Similarly to Eq.(23), Eq.(25) can further be approximated at $T\sim 300$ MeV to be
$V_{\rm eff}^{\rm(Tad)}(\alpha)\Bigg{|}_{\theta=0}^{T\sim 300\,{\rm MeV};eB={\cal O}(f_{\pi}^{2})}$ $\approx\frac{|eB|^{3}}{f_{0}}\frac{(2/3)^{2}}{8\pi^{2}}\frac{|\alpha-m_{0}|}{\alpha^{2}}{\cal I}(T/\alpha)$
$\propto\frac{|eB|^{3}}{f_{0}}\frac{|\alpha-m_{0}|}{\alpha^{2}}\,.$ (27)
This does not develop any peak in contrast to the $\theta=\pi$ case. Note that both Eq.(24) and Eq.(27) become identical when $m_{0}=0$. In other words, the discrepancy between two cases (with $\theta=0$ or $\pi$) has been generated by nonzero current quark mass.
Other remarks are in order:
  • the electromagnetic interaction does not break the CP symmetry, hence does not destabilize the symmetric phase $\alpha(\theta=\pi)\propto P=0$, which is reflected in the potential form in Eq.(24), while does explicitly break the chiral (and $U(1)$ axial) symmetry, therefore, the chiral symmetric phase $\alpha(\theta=0)\propto S=0$ is no longer stabilized, as in Eq.(27);
  • in the chiral limit $m_{0}\to 0$, the $\theta$ parameter becomes unphysical, so the CP order parameter $P$ coupled to $\theta$ as well as the CP phase transition is ill-defined unless another CP violating parameter is introduced. Therefore, the discrepancy between two cases (with $\theta=0$ or $\pi$) should be seen as long as quarks are massive only when the $\theta$ parameter is physical.
The phase transitions along the chiral $P^{\prime}(=S)$ and Polyakov loop ($\Phi$) directions are smooth crossover, where $P^{\prime}(=S)\sim 0$ at any $T$. Those crossover transitions proceed not much different from the case without the electromagnetic scale anomaly, nor $eB$ Sakai:2011gs ; Huang:2024nbd at $\theta=\pi$. See Fig. 4 showing the pseudocriticality of the deconfinement $(\Phi=1)$ confinement $(\Phi=0)$ transition, in which we observe $T_{\rm pc}^{({\rm dec})}\simeq 258$ MeV. Thus, the presence of the electromagnetic scale anomaly supports the generation of the CP-broken deconfinement(-like) domain $T^{\rm(CP)}_{c}>T^{\rm(dec)}_{\rm pc}$. This domain gets wider as $eB$ increases up to $\sim m_{\rm dyn}^{2}$.
Figure 4: The plot of the Polyakov loop variable $\Phi$ as a function of $T$, for the same parameter choice as in Fig. 1. All the curves look like overlapped including the case without the electromagnetic scale anomaly or even $eB$, which implies that the deconfinement phase transition is overall fairly insensitive to $eB$ in the present model study.
IV Summary and discussion
In summary, we have discussed the thermal CP phase transition in QCD at $\theta=\pi$ under a weak magnetic field background with $eB\lesssim m_{\rm dyn}^{2}$, where the electromagnetic scale anomaly plays a significant role. We have worked on a two-flavor PNJL model at $\theta=\pi$ in the MFA, including the electromagnetic-scale anomaly term (Eqs.(14) with (16)). The thermal CP phase transition has been observed to be of first order, in sharp contrast to the two-flavor (P)NJL without the electromagnetic scale anomaly (Figs. 1 and 2).
The electromagnetic-scale anomaly effect is dominated by the thermomagnetic term (the second term in Eq.(16)), which significantly creates the potential barrier at high temperatures around the origin of the CP-order parameter ($P$) direction (Fig. 3). This barrier generation is induced by non-standard potential or nonperturbative form $\sim|eB|^{3}\frac{|P|}{P^{2}+m_{0}^{2}}$ (Eq.(24)), in contrast to the CP-invariant chiral phase transition at $\theta=0$ (Eq.(27)), although both forms are reduced to be identical each other in the chiral limit $m_{0}\to 0$. This implies a significant amplification of the $U(1)$ axial violation due to the electromagnetic scale anomaly at high temperatures when $\theta=\pi$.
Since being insensitive to the electromagnetic scale anomaly, the phase transition along the Polyakov loop direction is not much different from the case without the electromagnetic scale anomaly, therefore goes like smooth crossover with the pseudocriticality. Thus, the presence of the electromagnetic scale anomaly supports the generation of the CP-broken deconfinement(-like) domain $T^{\rm(CP)}_{c}>T^{\rm(dec)}_{\rm pc}$, which becomes wider as $eB$ increases.
The present analysis has been based on the model of conventional NJL type, with which the inverse magnetic catalysis is not realized in a strong $eB$ regime at $\theta=0$, in contradiction to the lattice result Bali:2011qj . The electromagnetic scale anomaly survives only in a weak $eB$ regime, like $\sqrt{eB}\lesssim 100$ MeV, while the current lower bound of the magnetic field strength $\sqrt{eB}$ on lattices is $\sim$ 100 MeV Bali:2011qj . The inverse magnetic catalysis has therefore not yet been observed in such a weak $eB$ regime. Nevertheless, it would be intriguing to incorporate an $eB$-dependence into the scalar-four fermion coupling ($g_{s}$) to be fitted to the lattice data on the $eB$-dependence reflecting the inverse magnetic catalysis of the chiral crossover, as has been done in the literature even at $\theta=\pi$ Chatterjee:2014csa ; Bandyopadhyay:2019pml ; Carlomagno:2025ayh , to observe how the first-order nature of the CP phase transition could be weaker or still stay strong enough against the inverse magnetic catalysis. This is to be explored elsewhere.
An extended Columbia plot at $\theta=\pi$ with varying quark masses under a weak magnetic field background can also straightforwardly be investigated in a way similar to the work in Kawaguchi:2021nsa for the chiral phase transition. In the present analysis, we have also found that the potential barrier generation thermomagneticly induced from the electromagnetic scale-anomaly seems to persist as long as the up and down quark masses are in a range $1010^{-5}<m_{0}/m_{\rm phys}<$ 10, where $m_{\rm phys}$ denotes the value at the physical point. In the case below the lower bound, the CP-even and -odd chiral-order parameters $S$ and $P$ behave like almost identical $(V_{\rm eff}^{\rm(Tad)}\propto 1/|\alpha|)$ in the tadpole term contribution (see Eqs.(27) and (24)), hence the intrinsic barrier structure is gone in the $P$-direction. When the quark mass exceeds the upper bound, no spike structure or nontrivial contribution in the $P$-direction will be left. The parameter space in the two-flavor (P)NJL model outside the first-order regime above would all be the second order regime, unless another explicit CP violation other than what is induced from nonzero $\theta$ is introduced. The case would become nontrivial when the strange quark is incorporated into the model, because of the intrinsic potential barrier structure at $T=0$ associated with the cubic potential term arising from the $U(1)$ axial anomaly. Detailed study on this issue deserves another publication.
At any rate, the presently observed first-order nature of the CP phase transition will provide a new benchmark toward understanding of QCD at $\theta=\pi$, prior to the future development in lattice simulations and the ’t Hooft anomaly matching condition. Though being effective enough to make such a first step, the current model analysis can be improved by going beyond the MFA, say, using the functional renormalization group method.
In closing, we give comments on the phenomenological and cosmological applications. The first-order criticality of the CP phase transition in a weak magnetic background could give an impact on the interpretation of the observed nano-hertz gravitational-wave in terms of axionlike particle-domain wall collapses, as was addressed in Huang:2024nbd , and/or might also be relevant to the generation of primordial black hole. Those are also noteworthy to be pursued in another publication.
Acknowledgments
This work was supported in part by the National Science Foundation of China (NSFC) under Grant No.11747308, 11975108, 12047569, and the Seeds Funding of Jilin University (S.M.). The work by M.K. is supported by RFIS-NSFC under Grant No. W2433019. The work of A.T. was partially supported by JSPS KAKENHI Grant Numbers 20K14479, 22K03539, 22H05112, and 22H05111, and MEXT as “Program for Promoting Researches on the Supercomputer Fugaku” (Simulation for basic science: approaching the new quantum era; Grant Number JPMXP1020230411, and Search for physics beyond the standard model using large-scale lattice QCD simulation and development of AI technology toward next-generation lattice QCD; Grant Number JPMXP1020230409).